# Multi-Agent Inverted Transformer for Flight Trajectory Prediction

### v1

@@ I Introduction
Early models based on classical multivariate regression techniques have been proposed for flight trajectory prediction [7, 8]. Several studies have explored the use of Gaussian mixture models (GMMs) to represent the statistical distributions of flight trajectories and apply them to flight trajectory prediction [9, 10, 11, 12, 13].
Recent advancements have increasingly focused on deep learning-based approaches. Several models for flight trajectory prediction based on long short-term memory (LSTM) networks have been proposed [14, 15, 16]. Other studies have combined convolutional neural networks (CNNs) with LSTMs to improve the extraction of the spatio-temporal patterns of flight trajectories [17, 18].
@@
Building upon previous studies, this paper proposes a Multi-Agent Inverted Transformer (MAIFomer) for flight trajectory prediction. Aircraft movements in airspace are inherently multi-agent in nature, and an effective flight trajectory prediction model must account for both the spatio-temporal dynamics of individual aircraft and their social interactions with surrounding aircraft. The standard attention-based approach to modeling the interactions typically relies on computing attention scores between trajectory points in different time steps across different aircraft [24, 25]. Although this approach has demonstrated competitive performance in improving the prediction accuracy, it often yields attention distributions that are either irrelevant or difficult to interpret from a human perspective. In contrast, the proposed MAIFormer treats the entire trajectory of each aircraft as a single token and computes attention between agents (i.e., aircraft) to capture their interactions, as illustrated in Figure 1.
@@
The MAIFormer architecture has two distinct attention stages. In the first stage, masked multivariate attention (MMA) is used to capture the series-wise relationships among variates for each individual aircraft trajectory. In the second stage, agent attention (AA) is applied across different flights to model social interactions between aircraft. Unlike conventional approaches, MAIFormer decouples the processes of modeling intra-agent and inter-agent relationships into separate attention stages. This separation improves the accuracy of the prediction outputs and the interpretability of attention distributions. Interpretability is particularly important in safety-critical domains such as ATM, where the outputs of a flight trajectory prediction model must be understandable by human operators such as ATCs.
@@
  • The proposed model demonstrates significant performance improvements in prediction accuracy over other methods when evaluated on a real-world automatic dependent surveillance-broadcast (ADS-B) flight trajectory dataset, including the single-aircraft Transformer and AgentFormer.
  • MAIFormer uses a hierarchical attention mechanism to assign a single interpretable attention score to each aircraft, which enables a more intuitive understanding of how the model perceives inter-agent interactions within a traffic scene. This abstraction improves the model’s transparency and usability for human operators, such as ATCs.
@@ II-C Air Traffic Scene Tokenization
$\mathbf{C}=(\mathbf{c}^{1}_{1},\mathbf{c}^{2}_{1},\dots,\mathbf{c}^{F}_{1},\dots,\mathbf{c}^{1}_{N},\dots\mathbf{c}^{F}_{N})\in\mathbb{R}^{(N\cdot F)\times D}$ denotes the sequence of variate tokens, where each token is generated by applying a linear embedding to the original time series of each variate in $\mathbf{\overline{X}}$. Notably, the positional encoding is not applied after the tokenization, since there is no inherent ordering among agents in the air traffic scene.
@@ II-D MAIFormer: Multi-Agent Inverted Transformer
$\mathbf{C}_{\text{ST}}^{\ell-1}$ $=\text{MaskedMultivariateAttention}(\text{LN}(\mathbf{C}^{\ell-1})),$ (4)
@@
where $\mathbf{C}^{\ell}$ is the sequence of variate tokens from the $\ell$-th layer, and the output of the MMA module is $\mathbf{C}_{\text{ST}}$. The sequence of agent tokens $\mathbf{C}_{\text{A}}$ is then obtained by reshaping $\mathbf{C}_{\text{ST}}$ through agent-wise concatenation, and the output of the AA module is $\mathbf{C}_{\text{SC}}$. To improve the convergence and training stability, layer normalization (LN) was applied to the input of each module.
@@
2:  $\mathbf{C}_{\text{ST}}^{\ell-1}=\text{MMA}(\text{LN}(\mathbf{C}^{\ell-1}))$ $\triangleright\mathbf{C}_{\text{ST}}^{\ell-1}\in\mathbb{R}^{(N\cdot F)\times D}$
@@ II-D2 Agent Attention
where each $\mathbf{W}_{\text{SC}}\in\mathbb{R}^{(F\cdot D)\times(F\cdot D)}$ is a learnable projection for agent tokens that transforms the agent tokens into agent queries $\mathbf{Q}_{\text{SC}}$, agent keys $\mathbf{K}_{\text{SC}}$, and agent values $\mathbf{V}_{\text{SC}}$.
@@ III-A Dataset Description
To address the irregular sampling rate of ADS-B data, we resampled all trajectories at a uniform 6-second interval using the piecewise cubic Hermite interpolating polynomial (PCHIP) [29]. The trajectory dataset was then divided into training $\mathcal{D}_{\text{train}}$, validation $\mathcal{D}_{\text{val}}$, and test $\mathcal{D}_{\text{test}}$ sets using an 8:1:1 ratio (the first 80% of samples was used for training, the next 10% was used for validation, and the final 10% was used for testing).
For each data split, we constructed a past air traffic scene $\mathbf{X}$ and a future air traffic scene $\mathbf{Y}$ over 2-minute intervals (i.e., 20 time steps at 6-second intervals). This was done by sliding a window with one time step (6 seconds), as illustrated in Figure 5. To normalize the data, we applied min-max normalization to all variates using the minimum and maximum values derived from the training set. The normalization was independently applied for each variate to account for differences in their scales. As a result, we obtained a total of 509,389 traffic scenes for our experiments.
@@ III-C Comparison Baselines
We compared the downstream performance of MAIFormer for flight trajectory prediction with that of Transformer [27] and AgentFormer [23]. To ensure fair comparison, all models are configured with the same key hyperparameters as MAIFormer, including the latent dimension, number of layers, and feedforward network dimension. To compare the prediction accuracy, we used the mean absolute error (MAE), root mean squared error (RMSE), and mean absolute percentage error (MAPE). MAE, RMSE, and MAPE were applied separately to the latitude, longitude, and altitude dimensions. MAE provides the average magnitude of absolute prediction errors, while RMSE squares the errors before averaging and gives more weight to larger errors, which makes it sensitive to outliers. MAPE was used to evaluate relative error sizes across dimensions with different scales. These evaluation metrics are defined as follows:
@@ IV-A Quantitative Analysis
MAIFormer substantially outperformed all baseline models. The average MAE of MAIFormer shows the reductions of 22.46%, 16.93%, and 37.93% compared to the second-best model in the latitude, longitude, and altitude dimensions, respectively. Improvements in RMSE were even more pronounced, with error reductions of 33.33%, 18.14%, and 41.23% in the corresponding dimensions. We attribute these stronger RMSE gains to MAIFormer’s ability to more effectively capture the underlying dynamics of air traffic scenes. MAIFormer also achieved MAPE reductions of 23.01%, 15.27%, and 33.85% in latitude, longitude, and altitude dimensions, respectively.
Figure 6: Flight trajectory prediction results for two illustrative flight scenarios. Each row represents a different scenario, and each column shows the predictions from a different model.
When analyzing errors across different prediction horizons, the Transformer performed slightly better at early time steps (e.g., horizons 1 and 5) compared to AgentFormer and MAIFormer. The reason is likely that short-term future flight trajectories are primarily governed by the dynamics of individual aircraft, which makes them easier to extrapolate without the need to explicitly model agent interactions. In contrast, longer-term future trajectories (e.g., horizons 10, 15, and 20) are more influenced by surrounding aircraft and overall traffic patterns. Therefore, Transformer underperforms in longer-term predictions due to its inability to account for the context of neighboring traffics.
@@
IV-B Qualitative Analysis
For qualitative evaluation, we visualized the flight trajectory prediction results for two illustrative scenarios in Figure 6, which correspond to arrivals at ICN Runway 15, and Runway 33. In each plot, the blue line represents the past trajectory (used as input to the models), the green line indicates the ground-truth future trajectory, and the red line represents the model’s prediction.
For the Runway 15 arrival scenario (top row), MAIFormer accurately predicts the future trajectories of AC1, AC2, and AC3, even in the dense airspace near the runway. Both the Transformer and AgentFormer produce relatively accurate predictions for aircraft located farther from the runway (e.g., AC9 and AC10), but they struggle to maintain accuracy when predicting the trajectories of closely spaced aircraft near the runway.
For the Runway 33 arrival scenario (bottom row), both AgentFormer and MAIFormer demonstrate accurate overall predictions. However, AgentFormer fails to correctly capture the AC9’s future trajectory and incorrectly predicts that it will arrive ahead of AC8, which is already established on a structured arrival flow. In contrast, MAIFormer successfully captures the subtle interactions between the two aircraft and correctly places AC9 behind AC8 in the sequence. This example highlights MAIFormer’s better capability of understanding traffic scenes and inter-agent interaction modeling.

### v2

@@ I Introduction
Early models based on classical multivariate regression techniques have been proposed for flight trajectory prediction [6, 16]. Several studies have explored the use of Gaussian mixture models (GMMs) to represent the statistical distributions of flight trajectories and apply them to flight trajectory prediction [2, 23, 26, 17, 3]. Recent advancements have increasingly focused on deep learning-based approaches. Several models for flight trajectory prediction based on long short-term memory (LSTM) networks have been proposed [29, 39, 40]. Other studies have combined convolutional neural networks (CNNs) with LSTMs to improve the extraction of the spatio-temporal patterns of flight trajectories [22, 18]. Social-LSTM networks also have been proposed for multi-agent flight trajectory prediction [35, 5].
@@
Aircraft movements in airspace are inherently multi-agent in nature, and an effective model for flight trajectory prediction must account for both the spatio-temporal dynamics of individual aircraft and their social interactions with surrounding aircraft. Standard attention-based approaches typically rely on computing attention scores between trajectory points at different time steps for different aircraft [38, 4, 7]. These approaches handle spatio-temporal dynamics and social interactions jointly within a single attention layer, which inevitably increases modeling complexity and reduces the interpretability of the resulting attention scores. Several recent studies have explored hierarchical attention mechanisms to capture multi-level traffic information more efficiently, but these approaches still primarily rely on low-level time-step-to-time-step attention [37, 31, 34]. Consequently, the learned attention distributions often remain irrelevant or hard to interpret from a human perspective.
Motivated by these limitations, this paper proposes a Multi-Agent Inverted Transformer (MAIFormer) for flight trajectory prediction. MAIFormer reflects the hierarchical nature of information in multi-agent flight trajectories by explicitly decoupling the modeling of intra-agent and inter-agent relationships into two dedicated attention stages. In the first stage, masked multivariate attention (MMA) is used to capture the series-wise relationships among variates for each individual aircraft trajectory. In the second stage, agent attention (AA) is applied across different flights to model social interactions between aircraft. This hierarchical design separately represents intra-agent trajectory evolution (low-level) and inter-agent influence (high-level) and is better aligned with the reasoning process of human ATCs.
Furthermore, MAIFormer represents the entire trajectory of each aircraft as a single token and computes attention directly between agents (i.e., aircraft) to capture their high-level interactions, as illustrated in Figure 1. This agent-level attention mechanism leads to more structured and interpretable attention distributions. Interpretability is particularly important in safety-critical domains such as ATM, where the outputs of a flight trajectory prediction model must be understandable by human operators such as ATCs.
@@
  • The proposed model demonstrates significant performance improvements in prediction accuracy over other methods when evaluated on a real-world automatic dependent surveillance-broadcast (ADS-B) flight trajectory dataset.
  • MAIFormer uses a hierarchical attention mechanism to assign a single interpretable attention score to each aircraft, which enables a more intuitive understanding of how the model perceives inter-agent interactions within a traffic scene. The attention entropy analysis shows that MAIFormer yields lower attention entropy compared with AgentFormer, indicating more concentrated and thus more interpretable interaction patterns for human operators.
@@ II-C Air Traffic Scene Tokenization
$\mathbf{C}=(\mathbf{c}^{1}_{1},\mathbf{c}^{2}_{1},\dots,\mathbf{c}^{F}_{1},\dots,\mathbf{c}^{1}_{N},\dots\mathbf{c}^{F}_{N})\in\mathbb{R}^{(N\cdot F)\times D}$ denotes the sequence of variate tokens, where each token is generated by applying a linear embedding to the original time series of each variate in $\mathbf{\overline{X}}$. Notably, the positional encoding is not applied after the tokenization, since there is no inherent ordering among agents in the air traffic scene. As a result, the input to the MAIFormer consists of $N\cdot F$ tokens, which are each a $D$-dimensional representation vector corresponding to each aircraft variate.
@@ II-D MAIFormer: Multi-Agent Inverted Transformer
$\mathbf{C}_{\text{ST}}^{\ell-1}$ $=\text{MaskedMultivariateAttention}(\text{LN}(\mathbf{C}^{\ell-1}))+\mathbf{C}^{\ell-1},$ (4)
@@
where $\mathbf{C}^{\ell}\in\mathbb{R}^{(N\cdot F)\times D}$ is the sequence of variate tokens from the $\ell$-th layer, and the output of the MMA module is $\mathbf{C}_{\text{ST}}\in\mathbb{R}^{(N\cdot F)\times D}$. The sequence of agent tokens $\mathbf{C}_{\text{A}}\in\mathbb{R}^{N\times(F\cdot D)}$ is then obtained by reshaping $\mathbf{C}_{\text{ST}}$ through agent-wise concatenation, and the output of the AA module is $\mathbf{C}_{\text{SC}}\in\mathbb{R}^{(N\cdot F)\times D}$. To improve the convergence and training stability, layer normalization (LN) was applied to the input of each module.
@@
2:   $\mathbf{C}_{\text{ST}}^{\ell-1}=\text{MMA}(\text{LN}(\mathbf{C}^{\ell-1}))+\mathbf{C}^{\ell-1}$ $\triangleright\mathbf{C}_{\text{ST}}^{\ell-1}\in\mathbb{R}^{(N\cdot F)\times D}$
@@ II-D2 Agent Attention
where each $\mathbf{W}_{\text{SC}}\in\mathbb{R}^{(F\cdot D)\times(F\cdot D)}$ is a learnable projection for agent tokens that transforms the agent tokens into agent queries $\mathbf{Q}_{\text{SC}}\in\mathbb{R}^{N\times(F\cdot D)}$, agent keys $\mathbf{K}_{\text{SC}}\in\mathbb{R}^{N\times(F\cdot D)}$, and agent values $\mathbf{V}_{\text{SC}}\in\mathbb{R}^{N\times(F\cdot D)}$.
@@ III-A Dataset Description
To address the irregular sampling rate of ADS-B data, we resampled all trajectories at a uniform 6-second interval using the piecewise cubic Hermite interpolating polynomial (PCHIP) [11]. The resampling interval of 6 seconds was selected as a proxy for the average update rate of approximately 5 seconds in the original ADS-B data [10]. Finer temporal resolutions result in excessively long sequences with limited performance gains, while coarser resolutions fail to capture the short-term spatio-temporal dynamics of aircraft movement. The trajectory dataset was then divided into training $\mathcal{D}_{\text{train}}$, validation $\mathcal{D}_{\text{val}}$, and test $\mathcal{D}_{\text{test}}$ sets using an 8:1:1 ratio (the first 80% of samples was used for training, the next 10% was used for validation, and the final 10% was used for testing).
For each data split, we constructed a past air traffic scene $\mathbf{X}$ and a future air traffic scene $\mathbf{Y}$ over 2-minute intervals (i.e., 20 time steps at 6-second intervals). This was done by sliding a window with one time step (6 seconds), as illustrated in Figure 5. The prediction horizon of 2 minutes was motivated by operational considerations in air traffic control as it corresponds to the typical look-ahead time for short-term conflict alert (STCA) system for ATCs [9]. Nevertheless, the proposed method is not restricted to this horizon, and other time horizons can be readily applied and should be explored in future research. To normalize the data, we applied min-max normalization to all variates using the minimum and maximum values derived from the training set. The normalization was independently applied for each variate to account for differences in their scales. As a result, we obtained a total of 509,389 traffic scenes for our experiments.
@@ III-C Comparison Baselines
We compared the downstream performance of MAIFormer for flight trajectory prediction with that of Transformer [33], PatchTST [24], and FlightBERT++ [13] (single-agent methods for flight-trajectory prediction), as well as Social-LSTM [1] and AgentFormer [38] (multi-agent methods). To ensure fair comparison, all models are configured with the same key hyperparameters as MAIFormer, including the latent dimension, number of layers, and feedforward network dimension. To compare the prediction accuracy, we used the mean absolute error (MAE), root mean squared error (RMSE), and mean absolute percentage error (MAPE). MAE, RMSE, and MAPE were applied separately to the latitude, longitude, and altitude dimensions. MAE provides the average magnitude of absolute prediction errors, while RMSE squares the errors before averaging and gives more weight to larger errors, which makes it sensitive to outliers. MAPE was used to evaluate relative error sizes across dimensions with different scales. These evaluation metrics are defined as follows:
@@ IV-A Quantitative Analysis
Figure 6: Flight trajectory prediction results for two illustrative flight scenarios. Each scenario is presented in two rows showing predictions from six different models.
MAIFormer substantially outperformed all baseline models. Compared to the second-best model, the average MAE of MAIFormer showed reductions in the latitude, longitude, and altitude dimensions of 21.11%, 9.81%, and 34.88%, respectively. Improvements in RMSE were even more pronounced, with error reductions of 32.64%, 13.93%, and 38.07% in the corresponding dimensions. We attribute these stronger RMSE gains to MAIFormer’s improved ability to capture the underlying dynamics of air-traffic scenes. MAIFormer also achieved MAPE reductions of 16.67%, 3.66%, and 22.08% in latitude, longitude, and altitude dimensions, respectively.
When analyzing prediction errors for different prediction horizons, the single-agent models (Transformer, PatchTST, and FlightBERT++) slightly outperformed the multi-agent models Social-LSTM, AgentFormer, and MAIFormer at early time steps (e.g., horizons 1 and 5). This behavior is likely due to short-term future flight trajectories being largely dominated by the dynamics of individual aircraft, such as current velocity and heading, which means that they can be extrapolated easily without explicitly modeling interactions with surrounding traffic. In contrast, longer-term future trajectories (e.g., horizons 10, 15, and 20) are more strongly influenced by surrounding aircraft and overall traffic patterns. As a result, single-agent models generally underperform in longer-term predictions due to their inability to account for the contextual interactions between traffic. Notably, compared with other multi-agent models (Social-LSTM and AgentFormer), MAIFormer achieves more consistent and substantial performance gains over single-agent models across all prediction horizons.
IV-B Qualitative Analysis
For qualitative evaluation, we visualized the flight trajectory prediction results for two illustrative scenarios in Figure 6, which correspond to arrivals at ICN Runway 15, and Runway 33. In each plot, the blue line represents the past trajectory (used as input to the models), the green line indicates the ground-truth future trajectory, and the red line represents the model’s prediction. Notably, in the Runway 15 arrival scenario (top two rows of Figure 6), more complex and irregular traffic patterns occur near the airport compared with the Runway 33 scenario. This is due to the northern part of the ICN airspace being largely unavailable due to its proximity to the North Korean border, which requires arriving traffic to merge earlier than in the Runway 33 scenario.
For the Runway 15 scenario, single-agent models struggle more to accurately predict trajectories for aircraft near the runway compared to multi-agent models. In the vicinity of the runway, inter-aircraft separation becomes small, which makes capturing aircraft interactions crucial. The multi-agent models, Social-LSTM and AgentFormer, perform reasonably well even in the dense area near the runway, but their predicted trajectories for AC4 and AC5 are unrealistically close to each other. In contrast, MAIFormer consistently produces accurate predictions for all aircraft in this region, and its predicted trajectories maintain appropriate separation. For aircraft that are farther from the runway, both single-agent and multi-agent models generate relatively accurate predictions.
For the Runway 33 arrival scenario (bottom two rows of Figure 6), all baseline models perform reasonably well overall but still produce inaccurate predictions for several aircraft. For example, the baseline models fail to capture AC9’s future trajectory and incorrectly predict that it will arrive ahead of AC8, which is already established within a structured arrival flow. In contrast, MAIFormer successfully captures the subtle interactions between these two aircraft and correctly places AC9 behind AC8 in the arrival sequence. This example highlights MAIFormer’s stronger capability of understanding traffic scenes and modeling inter-agent interactions.
@@ IV-B Qualitative Analysis
IV-C Attention Entropy Analysis
We performed an attention entropy analysis to quantitatively evaluate and compare the interpretability of the agent-level attention patterns produced by MAIFormer and AgentFormer. For each query agent $i$, we computed the Shannon entropy of its attention distribution over all key aircraft $j$, which provides a scalar measure of how concentrated or diffused the learned interactions are. The attention entropy is defined as follows:
$H=\frac{1}{Q}\sum_{i=1}^{Q}\left(-\sum_{j=1}^{K}a_{ij}\log(a_{ij}+\epsilon)\right)$ (14)
where $Q$ and $K$ are the numbers of query and key tokens, respectively, $a_{ij}$ is the normalized attention weight assigned from query agent $i$ to key agent $j$, and $\epsilon$ is a small constant that was added for numerical stability. Lower entropy indicates that the attention is concentrated on a smaller subset of agents, while higher entropy indicates a more diffuse attention distribution. We report the attention entropy as a function of the number of aircraft in the traffic scene to examine how attention concentration varies with traffic complexity.
Figure 8: Comparison of attention entropy with respect to the number of aircraft.
TABLE II: Ablation studies. The best performance is highlighted in bold, and the second-best performance is underlined.
Model Horizon MAE $\downarrow$ RMSE $\downarrow$ MAPE (%) $\downarrow$
Lat () Lon () Alt (ft) Lat () Lon () Alt (ft) Lat Lon Alt
MAIFormer (w/o MMA) 1 0.0009 0.0015 33.2131 0.0010 0.0016 36.4879 0.0024 0.0012 0.4211
5 0.0018 0.0024 95.6454 0.0029 0.0033 132.4674 0.0041 0.0023 1.4441
10 0.0031 0.0040 155.4566 0.0040 0.0051 171.6697 0.0088 0.0035 2.7714
15 0.0039 0.0048 196.6767 0.0049 0.0056 282.4156 0.0095 0.0038 2.8912
20 0.0042 0.0056 225.7871 0.0052 0.0069 299.1411 0.0115 0.0047 3.4396
MAIFormer (w/o AA) 1 0.0009 0.0015 37.4491 0.0010 0.0017 43.1547 0.0025 0.0013 0.4227
5 0.0018 0.0026 119.0749 0.0028 0.0032 141.6576 0.0040 0.0022 1.4174
10 0.0029 0.0038 141.4881 0.0038 0.0048 166.6623 0.0085 0.0034 1.5551
15 0.0039 0.0047 201.7765 0.0048 0.0058 275.4671 0.0108 0.0039 2.8991
20 0.0043 0.0057 244.5641 0.0055 0.0070 301.1086 0.0116 0.0048 3.4912
MAIFormer (w/o inverting) 1 0.0011 0.0017 36.2344 0.0011 0.0018 45.1473 0.0028 0.0014 0.4474
5 0.0024 0.0030 142.1475 0.0032 0.0041 181.7415 0.0064 0.0031 2.3477
10 0.0036 0.0045 200.5464 0.0044 0.0058 264.1141 0.0121 0.0043 2.9991
15 0.0041 0.0052 221.4414 0.0052 0.0061 281.7987 0.0137 0.0045 3.0041
20 0.0049 0.0058 264.4875 0.0061 0.0074 322.1772 0.0145 0.0049 3.6468
MAIFormer (Proposed) 1 0.0008 0.0014 29.1527 0.0008 0.0016 33.5542 0.0023 0.0011 0.4104
5 0.0012 0.0020 60.1332 0.0014 0.0021 67.7342 0.0033 0.0015 0.8347
10 0.0022 0.0030 99.7254 0.0023 0.0034 115.0246 0.0056 0.0025 1.4943
15 0.0028 0.0041 131.7546 0.0033 0.0048 159.4441 0.0078 0.0034 2.1112
20 0.0037 0.0047 164.1826 0.0046 0.0065 198.2483 0.0101 0.0037 2.7827
As shown in Figure 8, as the number of aircraft grows, the attention entropy of AgentFormer increases substantially and remains relatively high even for small traffic scenes. In contrast, MAIFormer consistently exhibits lower attention entropy and a much slower growth rate with increasing traffic density, which indicates that the agent-level attention patterns are more concentrated and interpretable. Notably, the entropy curve of MAIFormer gradually converges, which suggests that even as the traffic density increases, MAIFormer does not diffuse its focus but instead attends to a limited set of key aircraft that most strongly influences the future trajectory of the query aircraft. Such focused attention patterns and stable entropy growth are desirable as they improve the usability of MAIFormer for interaction-aware decision support and analysis of interaction priorities in historical traffic data.
IV-D Ablation Studies
We also conducted ablation studies to examine the contributions of key components in MAIFormer. We removed the MMA module and allowed the AA module to capture both spatio-temporal dynamics and inter-agent interactions to assess whether aggregated agent tokens alone are sufficient for accurate prediction. We then retained MMA but removed AA and the masking strategy in MMA to evaluate how agent attention over variate tokens affects performance. Finally, we replaced the inverted embedding with a standard temporal embedding while keeping the rest of the MAIFormer architecture unchanged to quantify the impact of the embedding strategy.
Table II summarizes the ablation results. All ablated variants exhibit poorer performance compared to the full MAIFormer architecture. In particular, when either MMA or AA is removed, the model is still able to produce reasonable predictions but consistently shows greater prediction errors than the complete model. This suggests that while MMA or AA alone can partially capture relevant information, separating spatio-temporal dynamics and inter-agent interactions into dedicated hierarchical attention modules leads to more effective representation learning and better performance. Moreover, replacing the inverted embedding with standard temporal embedding results in the largest performance degradation. This finding indicates that conventional temporal embeddings are insufficient to capture complex and airspace-specific flight patterns, and the combination of the inverted embedding and hierarchical attention is essential for achieving the best prediction accuracy with MAIFormer.